(Handwritten step by step please) Please find the call and put option prices using the information below. You must show your work to receive credit.

Stock Price=45

Exercise Price=45

Risk-free rate=5%

Time to maturity=6 months

Standard deviation=30%
Answer #1

Risk Free Rate Factor (R)=1+5%=1.05

Upward Price factor(u)=1+30%=1.3

Downward price factor(d)=1-30%=0.7

Probability of going upward (p)= (R-d)/(u-d)=(1.05-0.7)/(1.3-0.7)=0.583

Probability of going downward(1-p)=1-0.583=0.417

Price of call option= (stock price*u-exercise price)*p/R=(45*1.3-45)*0.583/1.05=$7.49

Similarly, price of put option= (Exercise price-stock price*d)*(1-p)/R=(45-45*0.7)*0.417/1.05=$5.36